OpenStudy (anonymous):

Solving quadratic equations by factoring Find the solution set of x^2 + 4x - 5 = 0

OpenStudy (anonymous):

x^2+4x-5=(x+5)(x-1) And you get the solutions

OpenStudy (anonymous):

(x-1)(x+5)=0 which means x=1 or x=-5

OpenStudy (anonymous):

ye

OpenStudy (anonymous):

Oh, thanks for your help! I figured this one out a little while ago.